Two lines L1 :2y-3x-6=0and L2 :3y+x-20=0 intersect at point A. A third line L3 is perpendicular to L2 at point A. Find the equation of L3 in the Form y=mx+c where m and c are constants
Answer by Boreal(15235)

find the intersection point.
From L2, x=20-3y
so 2y-60+9y-6=0 by substitution
11y=66
y=6
x=2
They intersect at (2, 6)
L2 equation is also y=(-1/3)x+20/3
slope of the desired line is negative reciprocal of -1/3 or 3
point slope formula is y-y1=m (x-x1),m slope,(x1, y1) point
y-6=3(x-2)
y=3x or y=3x+0.